OpenStudy (anonymous):

What should be added to both sides of the equation to solve for the variable. A+98=-2

OpenStudy (lgbasallote):

hmmm remember that you're looking for A so all constants with A must become zero...what do you add to both sides so +98 will become 0?

OpenStudy (anonymous):

-2?

OpenStudy (lgbasallote):

is 98 - 2 = 0?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

idk...

OpenStudy (anonymous):

try to solve as @lgbasallote say

OpenStudy (anonymous):

-100?

OpenStudy (anonymous):

A+98=-2 ise A+98+(-98)=-2+(-98) ise A=-100 YOU ADD -98 to both side of the equality

OpenStudy (anonymous):

THanks